OpenStudy (anonymous):

what the y intercept and x intercept of y=2x-5

OpenStudy (anonymous):

x-intercept: y=2x-5 0=2x-5 5=2x x=5/2

OpenStudy (anonymous):

y-intercept- y=2x-5 y=2(0)-5 y=0-5 y=-5

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D2x-5